Circular No. 6597

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

              URL http://cfa-www.harvard.edu/cfa/ps/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

SAX J1750.8–2900 A. Bazzano, Istituto di Astrofisica Spaziale (IAS), CNR, Frascati; J.

Heise, Space Research Organization of the Netherlands (SRON), Utrecht; P. Ubertini, M. Cocchi, and L. Natalucci, IAS; J. M. Muller, SRON and BeppoSAX, Rome; R. Jager and J. in ’t Zand, SRON; and L. Di Ciolo, G. Celidonio, A. Coletta, P. Giommi, D. Ricci, M. Capalbi, and S. Rebecchi, BeppoSAX, Rome, report the detection of a new weak x-ray source in the galactic-center region with the Wide Field Camera onboard the BeppoSAX satellite: “The source, SAX J1750.8–2900, was initially detected on Mar. 18.032 UT, in the range 1.6–8.3 keV with an intensity of ∼ 50 mCrab, at α = 17^{h}50^{m}48^{s}, δ = −29^{o}00′.0 (equinox 2000.0; preliminary error radius 5′). The nearest known x-ray source is the ROSAT Point Source Catalogue object 1RXP J175029–2859.9 (flux 1.75 × 10^{−13} erg s^{−1} cm^{−2}), located 4′ away, at the edge of the SAX J1750.8–2900 error box. No optical, ultraviolet, or γ-ray sources are known within a 5′ radius. At radio wavelengths, three sources are listed in the NRAO/VLA Sky Survey database

(cf. http://www.cv.nrao.edu/~jcondon/nvss.html) at α = 17^{h}50^{m}29^{s}.9,

δ = −28^{o}59′07′′; 17^{h}50^{m}50^{s}.9, −29^{o}03′11′′; and 17^{h}51^{m}03^{s}.6, −29^{o}00′37′′. In the infrared, the IRAS 17477–2900 (from the Point-Source Catalogue) is nearby. During follow-up monitoring, a burst was detected on Mar. 20.38795 from the same position of the newly detected source, with a peak intensity of 0.7 Crab (1-s time bin) and duration 20 s.”

COMET 81P/WILD 2 T. Farnham and D. Schleicher, Lowell Observatory, report on narrow-

band photometry of comet 81P with the 0.79-m telescope on Mar. 5.4 UT, at r = 1.70: “Haser-model production rates were log Q(OH) = 27.77, log Q(CN) = 25.25, and log Q(C_{2}) = 24.92, and the dust-production rate was log (Afρ) = 2.63; combining these results with those previously obtained with the Hall telescope on Feb. 12.2 and 15.2 (r = 1.78–1.77 AU) implies that gas and dust production are essentially unchanged over this three-week interval.”

Further total visual magnitude estimates: Mar. 4.78 UT, 9.5 (K. Hornoch, Lelekovice, Czech Rep., 0.35-m reflector); 7.76, 9.6 (V. Znojil, Brno, Czech Rep., 25×100 binoculars); 11.92, 9.9 (A. Pereira, Cabo da Roca, Portugal, 0.25-m reflector); 13.09, 9.7 (J. E. Bortle, Stormville, NY, 0.41-m reflector); 16.91, 9.7 (B. H. Granslo, Fjellhamar, Norway, 0.20-m reflector).

1997 March 20  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Daniel W. E. Green